QUESTION 5 1 points Save Answer A triprotic acid, H3A has the following Kas: Ka1 = 4.22 x 10-4, Ka2 = 7.51 x 10-8 and Ka3 = 7A triprotic acid, H3A has the following Kas: Ka1 = 4.22 x 10-4 , Ka2 = 7.51 x 10-8 and Ka3 = 7.51 x 10-15. What is the Kb of H2A-?
Answer #1

Kb for HA2- = Kw/Ka2

Kb for HA2- = 10-14/7.51*10-8

Kb for HA2- = 1.33*10-7

Therefore, 1.33*10-7 is the correct answer.

Please give good rating.